OpenStudy (anonymous):

Find the radius of a cone when the surface area is 6pi in^2 and the height is 4".

OpenStudy (mayankdevnani):

right cone...yes or no

OpenStudy (mayankdevnani):

@tymiller

OpenStudy (mayankdevnani):

Surface = Pi*r^2 + Pi*r* sqrt (r^2+s^2)

OpenStudy (anonymous):

it doesnt say it just says conical paper cup

OpenStudy (anonymous):

here is the exact wording of the problem: A conical paper cup is to have a height of 4 inches. Find the radius r of the cone that will result in a surface area of 6π in2.

OpenStudy (anonymous):

The surface area is pi*radius*height. Plug in the numbers.

OpenStudy (amistre64):

conical paper sup; does the surface area include the inside of the cup as well?

OpenStudy (amistre64):



OpenStudy (amistre64):

http://math.about.com/od/formulas/ss/surfaceareavol_2.htm surface area is\[\pi r (s+r)\] lateral area is\[\pi rs\]

OpenStudy (anonymous):

Strictly they should also have said slant height (s) but I think its reasonable to assume that's what they meant here (else it's a bit messy).